OpenStudy (anonymous):

what is csc(1560°) and how do you solve it? I tried to put it in my calculator but I am getting the incorrect answer.

OpenStudy (anonymous):

csc is the inverse of sin. So when you put it in your calculator its a little different. try this (1/sin(1560))inverse. and then enter

OpenStudy (anonymous):

fisrt we need to find sin(1560) as csc(1560°) is the reciprocal of sin(1560)

OpenStudy (anonymous):

now sin(1560)=sin(4*360+ 120) =sin(120) =sin(180-60) ( u can write sin(120)=sin(90+30)=cos30) =sin60=sqrt(3)/2 hence csc(1560°) =1/sin(1560) =2/sqrt(3)

OpenStudy (anonymous):

Another approach:\[\text{Csc}[\text{Mod}[1560{}^{\circ},360{}^{\circ}]]=\frac{2}{\sqrt{3}} \]

OpenStudy (anonymous):

Thank you so much, I was extremely confused , the explanations you both gave were extremely helpful.